OpenStudy (andijo76):

=√ (200)(0.23)(0.77) how do i figure out the answer to this its part of my answer i get so far and when i hit this part i get stuck |

OpenStudy (andijo76):



OpenStudy (andijo76):

that is what it actually looks like

OpenStudy (ranga):

You multiply the three numbers and then take the square root.

OpenStudy (ranga):

200 x 0.23 x 0.77 = 35.42 sqrt(35.42) = 5.95

OpenStudy (andijo76):

oh ok thank you i kept getting stuck on this lol

OpenStudy (ranga):

you are welcome.

OpenStudy (phi):

notice you can write these numbers as 2*10^2 * 23*10^-2 * 7*11*10^-2 or 2*7*11*23 * 10^-2 sqr of 10^-2 is 10^-1 you get sqr(2*7*11*23)/10 notice that there is no perfect squares (all factors are prime)

OpenStudy (ranga):

Do they tell you if they want the answer in decimal or scientific notation or fraction or with radical?

OpenStudy (andijo76):

that is only part of the question but i did not figure it out right the question is a candy company claims that 13%of its plain candies are orange and a sample of 100 candies is randomly selected

OpenStudy (andijo76):

find the mean and standard deviation for the number of orange candies in such groups of 100

OpenStudy (andijo76):

okay i just confused you didnt i this is the same question but a different numbers i am not doing a test but a digital assignment i got a similar excersise

OpenStudy (ranga):

Mean = n * p where n is the sample size and p is the probability Here n = 100, p = 0.13 Mean = 100 * .13 = 13 Standard Deviation = sqrt(n * p * (1-p)) = sqrt( 100 * .13 * (1 - .13)) = sqrt( 11.31) = 3.36

OpenStudy (andijo76):

n represents 100 p represents 0.13 how do i find the value for q

OpenStudy (ranga):

q = 1- p

OpenStudy (andijo76):

according to the sample its a number i will show you the example

OpenStudy (andijo76):

r= represents the number of trials which is equal to 150 p= represents the probability of sucess in one of the trials which is equal to 0.18 q= represents the probability of a failure in one of the trials which is 0.82.

OpenStudy (andijo76):

i cant figure out how they got the 0.82.

OpenStudy (anonymous):

q=1-p=1-0.18 = 0.82

OpenStudy (andijo76):

oh ok

OpenStudy (andijo76):

i think this math is new to me i will figure it out i hope lol

OpenStudy (anonymous):

You will figure it out. Binomial distribution quite useful when something has only two possible outcomes (true/false, present/absent, etc.).